EVA Report – March 22nd

EVA Report

EVA #12 – 22th March 2018 – Sol 11

Crew members: Bastien Baix (EVA leader), Michael Saint-Guillain, Mario Sundic, Sophie Wuyckens

Objectives:
– Drone mapping and drone filming by Bastien
– Photos by Mario
– Sophie will put her detector outside of the RAM

Operations:
9:43 – Entering the airlock
9:47 – Leaving the MDRS with 2 rovers
9:55 – Arriving at Zubrin’s Head
10:20 – We start climbing a high hill
10:42 – Top of the mountain
10:50 – Bastien map the mountain
11:05 – We go down of the mountain
11:40 – We leave Zubrin’s Head
11:50 – Arrival at the MDRS, Mario and Bastien go inside, Michael and Sophie install the muon detector 12:05 – Michael and Sophie enter the airlock

Narrative:
One of our last EVA took place on soil 12 at 9:40. I am Bastien Baix and I was EVA leader today. We decided to climb a high hill today. To do so, we went to Zubrin’s Head. Near this location, we located our objective and we start our climbing. It was particularly difficult as the slope was nearly vertical! As Sophie is relatively clumsy, we made her a safety line…

This climbing allowed us to take amazing pics and vids with the drone. Once we made it to the top, the view was amazing. We’ve seen miles around.

When we decided to go back and reach our rovers, we couldn’t find them anymore! We pass half an hour to finally find our way back to them and ride to the MDRS. It was an awesome but sweaty EVA!

Vehicles:
Curiosity and Opportunity rovers

Operations Report – March 22nd

SOL: 11

Name of person filing report: Bastien BAIX
Non-nominal systems: noisy pump, see end of report below
Notes on non-nominal systems: –

Generator (hours run): Turned off at 12.40 pm and turned on at 6.30 pm Solar – SOC 80% (Before generator is run at night)
Diesel – ~25%
Propane – 80%
Ethanol Free Gasoline (5 Gallon containers for ATV) – 1 ½ full jerrycans.

Water (trailer) – 0 gallons.
Water (static) – ~500 gallons
Water (third tank) – 0 gallons
Trailer to Static Pump used – NO
Water (loft) – Static to Loft Pump used – YES
Water Meter: 132 585.7
Toilet tank emptied: NO

ATV’s Used: –
Oil Added: NO
ATV Fuel Used: NO
Hours the ATVs were used today: _
Notes on ATVs: –

Deimos rover used: NO
Hours: 116.7
Beginning charge: 100%
Ending charge: –
Currently charging: YES

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: NO
Hours: 26.9
Beginning charge: 100%
Ending charge: –
Currently charging: NO

Opportunity rover used: YES
Hours: 26
Beginning charge: 100%
Ending charge: 93%
Currently charging: NO

Curiosity rover used: YES
Hours: 20.8
Beginning charge: 100%
Ending charge: 91%
Currently charging: YES

HabCar used and why, where: –
General notes and comments: –
Summary of internet: ~ 120 Mb remaining
Summary of suits and radios: –
Summary of Hab operations: –
Summary of GreenHab operations: –
Summary of ScienceDome operations: –
Summary of RAMM operations: –
Questions, concerns and requests to Mission Support:

Update on the pump: I made pictures of the pump (and the old pump too). – Model: Pentair Shurflo
– 115 VAC
– 60Hz
– 3,3 GPM /12,5 LPM of max flow
– 104 Watts
– Pressure: 45 PSI / 3,1 bar

Sol Summary – March 22nd

Crew 190 Sol Summary Report
22 March 2018

Sol 11

Summary Title:
Happy Birthday!

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation for the last day on Mars (Sol 12), 08:00

Sol Activity Summary:
8:20 Gym
8:35 Medic inspection
8:50 Breakfast
9:40 EVA: Bastien (cartography), Sophie (placement of muon detector), Mario (photos), Michael (support) 12:05 End of EVA – Crew meeting – Relaxation
13:00 Lunch
14:00 Scientific & academic work – Interviews continued (Ariane) 18:00 Report redaction
19:00 CapCom

Look Ahead Plan:
Sol 12:
Notes: Last EVA for Maximilien & Ariane (soil sampling), Fred (collection of the experiment he placed outside several days ago), Martin (support). Part of the crew will conclude their experiments in the afternoon while the others will do some clean up inside the MDRS. Sol 12 will be the last day under simulation. On Sol 13 we will welcome Crew 191.

Anomalies in work:
None.

Weather:
Very cloudy all day long… Preventing from sun telescope activities, despite the sun storm announced. Too bad. Hopefully tomorrow will still allow us to observe residual sun activities.

Crew Physical Status:
Still tired, but we observe an improvement compared to the couple of last days! However we have significantly more difficulty to recover after the morning gym.

EVA:
See EVA report.

Reports to be filed:
* Commander report
* Journalist report
* EVA#12 report
* EVA#13 request for Sol12
* Engineer report
* Pictures
* Summary report

Support Requested:
None

Commander Report – March 22nd

Crew 190 – sol11 – crew commander report 22-03-2018

First, we made a short gym session for waking up. Then, we ate breakfast and we wished a happy birthday to Michael who is now 31 years old. Sophie, Bastien, Mario and Michael left for an EVA up to “Zubrin Head” place. They took some pictures and Bastien mapped the area. Afterwards, they installed the muons detector outside of the station before coming back inside the Hab.

At midday, Ariane started to cook the lunch and we spent time eating and speaking for Michael Birthday’s. To complete our afternoon, everyone worked to his scientific project. Last manipulations will be done tomorrow morning.
Tonight we will have dinner followed by a cake kindly prepared by the crew commander especially for Mick’s birthday.

Tomorrow, we plan to end the experiments and have a small EVA before starting to clean up the whole station. Our way back to Earth is now in the mind of each crew members.

Regarding the station, nothing to declare.

Journalist Report- March 22nd

This morning’s EVA (Extra Vehicular Activity) was a bit of a special one. Four of our crew members exited the vault, armed with a camera and a drone, and entered the 4×4 electrical rovers. They drove them south and stopped a bit further than the formation known as Zubrin’s Head. There, they exited the vehicles and walked west, towards the base of a single, conical, towering red mountain. At around 100m (300ft), it stood out as the tallest rock formation of the area, and they were ready to climb it. Despite sharply inclined slopes on either side, and protruding rocks near the top, the team was keen on making it to the top. With the drone, they circled the mountain, and determined an optimal path that went around the back side. On they went. This wasn’t the worst climb any of them had seen, but with the heavy space-suits on their backs, restricted vision, lack of water and limited supply of oxygen, it sure felt like one of the bad ones. Despite the difficulties, after half an hour of huff ‘n puff they finally reached the top, from which they could see the whole valley: snowy mountains far in the south, the Skyline Rim in the west, and the Martian plains stretching all the way to the horizon in the north, well past the MDRS and the area we have had the luck to explore these past two weeks.

This was their last EVA. On Saturday, they will embark on the shuttle which will take them away from this red land, away from the dry sand, the storms and the burning sun, and take them back home, back to Earth. Seeing, from above, the area they have spent these last two weeks in was a touching moment for them, and they felt the incoming goodbyes.

Back at base though, there were still things to do before departure. For starters, today was Michael’s birthday, and the whole team celebrated with birthday cake, board games and a custom-made paper tiara for Michael. The rest of the day was devoted to finishing up our experiments, and slowly beginning to clean up the base before the incoming crew gets here on Saturday. Indeed, there is only one day left for us here. And we’re trying to enjoy it as much as we can!

Greenhab Report – March 22nd

GreenHab report for Sol11

Crewmember’s name: Frédéric Peyrusson (Biologist)
Date 22/03/2018 (Mick’s birthday)

Environmental control:

Ambient with door open.
Shade cloth on.
Fan on, without cooling.

Average temperatures: 26°C
Low: 8°C
High: 32°C

Hours of supplemental light: 0 hours

Changes to crops: No change

Daily water usage for crops: 10 gal

Time(s) of watering for crops: 4

Moringa research observations: N/A

Changes to research plants: N/A

Mario finished his 3 hydroponic towers and transplanted lettuce and arugula. All other plants have been watered.

Support/supplies needed: N/A